### v1

# Analysis of Broken Randomized Experiments by Principal Stratification

Abstract
Although randomized controlled trials have long been regarded as the “gold standard” for evaluating treatment effects, there is no natural prevention from post-treatment events. For example, non-compliance makes the actual treatment different from the assigned treatment, truncation-by-death renders the outcome undefined or ill-defined, and missingness prevents the outcomes from being measured. In this paper, we develop a statistical analysis framework using principal stratification to investigate the treatment effect in broken randomized experiments. The average treatment effect in compliers and always-survivors is adopted as the target causal estimand. We establish the asymptotic property for the estimator. We apply the framework to study the effect of training on earnings in the Job Corps Study and find that the training program does not have an effect on employment but possibly have an effect on improving the earnings after employment.
Keywords: Broken randomized experiment; Causal inference; Non-compliance; Principal stratification; Truncation by death; Missing data.
1 Introduction
The randomized controlled trials (RCTs) have long been regarded as the “gold standard” for evaluating treatment effects, because randomization leads to identical distribution of baseline characteristics in the treated group and control group. Measured and unmeasured factors that affect the outcome are naturally balanced. As a result, the difference in average outcomes between the treated and control group must result from the treatment. However, although the effect of pre-treatment factors is eliminated using the randomized design, there is no guarantee that post-treatment events which can affect the primary outcome can be prevented. The distributions of baseline characteristics would be different if we only compare the individuals with measured and well defined outcomes.
There are some sorts of post-treatment events that lead to broken randomized experiments (Rubin, , 1976; Barnard et al., , 1998; Hirano et al., , 2000; Rubin, , 2006; McConnell et al., , 2008).
  1. 1.
    Non-compliance. The individuals do not adhere to the treatment assignments. For example, when studying the effect of a drug on a disease, some patients who were assigned to receive treatment are allergic to the treatment. So the practitioner offers these patients placebo even though the randomization computer assigns these patients to receive the active treatment. These patients the assigned treated group should be considered as the control when evaluating the treatment effect. However, the distributions of baseline characteristics are different between the actually treated individuals and actually controlled individuals, because there are more allergic individuals in the actually controlled group (Frangakis and Rubin, , 1999).
  2. 2.
    Truncation-by-death. The outcomes for some patients are not well defined. For example, to study the effect of a surgery on the quality of life, some patients die before the end of study. It is not reasonable to record the quality of life as any value because we can not define the quality of life for dead people. The observed survivors between the treated group and control group have different characteristics if the risks of mortality are different between groups. A popular terminology is the “survivor bias” (Zhou et al., , 2005; Snyder et al., , 2009). Another example is rescue medication, in which case the patients receive additional medication besides the assigned treatment if emergency events occur. The outcome would be resulted by the joint effect of the initial treatment and rescue medication. To evaluate the pure effect of the initial treatment, we need to separate the effect of rescue mediation from the joint effect.
  3. 3.
    Missingness. The outcome of interest is well defined, but is not collected. Missingness is different from truncation-by-death. For example, a patient quits the experiment because of moving to another city. Although this patient has an outcome (quality of life), the outcome is not recorded by the researcher.
Non-compliance, truncation-by-death and missingness can happen together in an experiment. We give two concrete examples to depict the broken randomized experiments.
Example 1 (Natural experiments)
Suppose the research interest is the return to an additional year of schooling. In economics, the season when a person was born is used as an instrumental variable. Those who were born late in a year tend to receive shorter education and enter the job market when they celebrate their adult birthday. It seems that the birth season is randomized by nature. However, some individuals would always drop out regardless of birth season, while some would always complete high school, resulting in non-compliance. Another problem is that if people are not employed, their earnings cannot be defined, resulting in truncation-by-death. Those staying at home may have excellent potential to earn money, and they just decide not to work because the salary offered by the employee does not meet their expectation. It is inappropriate to record his wage as zero.
Example 2 (Encouragement designs)
Suppose the research interest is the effect of vaccine on virus load after infection. Subject to ethical issues, practitioners can not force patients to take vaccine or not to take vaccine. Therefore, encouragement designs are usually adopted, where patients are randomly encouraged to take vaccine. With encouragement, patients are more likely to take vaccine. Those who are not encouraged can actually take vaccine, and those who are encouraged can reject vaccine by their one judgements, resulting in non-compliance. On the other hand, if a patient is not infected due to cautious self protection, the virus load cannot be defined, resulting in truncation-by-death. The vaccine may have no effect on infection but have significant effect on lowering the virus load, so we need measurements of virus load in the target population to complete the analysis.
Our work is motivated by the Job Corps (JC) education and job training program for disadvantaged youths (Schochet, , 2001; Burghardt et al., , 2001). In the Job Corps study, the disadvantaged youths are randomly provided with an opportunity to take a job training program, which lasts for several years. Employment status and weekly earnings were recorded. A lot of work has been done to study the effect of this job training program (Schochet et al., , 2008; Zhang et al., , 2009; Flores-Lagunes et al., , 2010; Flores et al., , 2012; Chen and Flores, , 2015). Nevertheless, the simultaneous existence of non-compliance and truncation-by-death provides great challenges to evaluate the pure effect of the program. Statistically sound and interpretable methods are still desired to analysis broken randomized experiments.
Methodologies have been proposed to deal with non-compliance. The most intuitive idea is to adopt the intention-to-treat strategy to study the effect of the treatment assignment (Ten Have et al., , 2008; McNamee, , 2009; Gupta, , 2011). The as-treated and per-protocol strategies focus on some subpopulations defined by observed compliance types. However, these estimands does not necessarily have meaningful causal interpretations, because the target population is defined by post-treatment variables which are essentially realizations of potential outcomes. The instrumental variable approach regards the treatment assignment as an instrumental variable for the actually received treatment (Angrist et al., , 1996; Martens et al., , 2006; Tan, , 2006; Baiocchi et al., , 2014). An homogeneity or non-interaction assumption is required to identify the treatment effect in the overall population (Wang and Tchetgen Tchetgen, , 2018).
The instrumental variable approach is closely related to principal stratification (Frangakis and Rubin, , 2002; Lipkovich et al., , 2022). With truncation-by-death, a principal stratum of always-survivors is adopted as the target population, among which both potential outcomes are well defined (Ding et al., , 2011). In economic studies with earnings as the primary outcome and employment as the intermediate variable, this principal stratum is termed the always-employed (Zhang and Rubin, , 2003; Zhang et al., , 2009). A monotonicity assumption on survival and a principal ignorability assumption (also referred to as explainable nonrandom survival) are usually made to identify the principal stratum treatment effect (Jo and Stuart, , 2009; Stuart and Jo, , 2015; Feller et al., , 2017; Jiang et al., , 2022; Qu et al., , 2023). Alternatives to principal ignorability include constructing bounds (Imai, , 2008; Long and Hudgens, , 2013; Yang and Small, , 2016), imposing selection models (Shepherd et al., , 2006; Jemiai et al., , 2007) and invoking substitutional variables (Tchetgen Tchetgen, , 2014; Wang et al., , 2017).
Truncation-by-death is a different problem with missingness. The former means that the outcome is not well defined, whereas the latter means that the outcome exists but is not observed. To deal with missing data, the selection model and pattern mixture model assume different data generating mechanisms and adopt different estimands (Little, , 1993, 2008). Usually, we do not incorporate the pattern of missingness in the principal strata. Under a missingness-at-random assumption, the target estimand can be estimated by likelihood methods (Little et al., , 2017).
Mattei and Mealli, (2007) extended the framework of principal stratification to randomized experiments suffering from treatment non-compliance, missing outcomes and truncation by death. In their work, there are only two types of compliance behaviors, namely compliers and never-takers. They developed a Bayesian model to estimate the causal parameters. Frumento et al., (2012) assumed a mixture distribution for potential outcomes and identified the parameters using EM algorithm. As a weakness of mixture models, idenitification is fragile when the outcome behaviors are similar between principal strata. Chen and Flores, (2015) constructed a bound for the average treatment effect in a principal stratum. The validity of bound relies on other assumptions and can be wide, so it may be inconvenient for inference based on the bound. There is a lack of theoretical properties on the estimators, which hinders statistical inference about related scientific questions.
The remainder of this article is organized as follows. Section 2 introduces the notations and assumptions for identification. We provide the identification results in the prensence of noncompliance, truncation-by-death and missingness. Section 3 introduces the estimation and inference procedure. We establish the asymptotic property for the estimator of principal stratum average treatment effect. In Section 4, we conduct simulation studies to assess the performance of the treatment effect estimator and the coverage of the confidence interval obtained from the asymptotic formula. In Section 5, the proposed framework is applied to study the effect of a job training program on the earnings in the principal stratum comprised of those who would always comply to the treatment assignment and always be employed regardless of training. Finally we point out some limitations and research directions in Section 6.
2 Framework, estimand and identification
2.1 Notations and assumptions
Let $Z$ be the binary treatment assignment, where $Z=1$ denotes the active treatment and $Z=0$ denotes the control. Following the potential outcomes framework (Rubin, , 1974), let $D(z)$ be the binary treatment an individual actually received if it is assigned to take the treatment $z$. Let $S(z,d)$ be the binary indicator of whether the outcome of interest is defined. In the truncation-by-death context, $S(z,d)=1$ means the individual survives and $S(z,d)=0$ means the individual dies under the treatment assignment $z$ and actual treatment $d$. In the employment context, $S(z,d)=1$ means employed and $S(z,d)=0$ means unemployed. Let $Y(z,d)$ be the potential outcome of interest, which is only well defined when $S(z,d)=1$. We supplementarily denote $Y(z,d)=*$ if $S(z,d)=0$.
We make the stable unit treatment value assumption (SUTVA) and consistency. That is, there is no interference between individuals. The potential variables $S(z,d)$, $Y(z,d)$ can be observed, denoted by $S$, $Y$ and $\Delta$, respectively, if $Z=z$ and $D=d$ in the experiment. Let $\Delta_{S}$ be the missingness indicator of the survival variable ($S$) and $\Delta_{Y}$ be the missingness indicator of the primary outcome ($Y$), with 1 indicating observation and 0 missingness. Suppose the sample includes $n$ independent individuals randomly drawn from a super-population. We use the subscript $i$ to denote the measurements for the $i$th individual when necessary. The observed sample is $\mathcal{O}=\{O_{i}=(Z_{i},D_{i},\Delta_{Si},S_{i}\Delta_{Si},\Delta_{Yi},Y_{i}\Delta_{Yi}):i=1,\ldots,n\}$.
We make the following assumptions, which are common in literature on instrumental variables (non-compliance).
Assumption 1 (Randomization)
[…]
Assumption 2 (Monotonicity on compliance)
$D(1)\geq D(0)$ almost surely.
Assumption 3 (Exclusion restriction)
$S(1,d)=S(0,d)$ for $d=0,1$. $Y(1,d)=Y(0,d)$ if $S(1,d)=S(0,d)=1$ for $d=0,1$.
Assumption 4 (Substitution relevance)
$E\{D(1)\}\neq E\{D(0)\}$.
Assumption 1 states that the treatment assignment is completely randomized. There should be no measured and unmeasured confounding. Therefore, the treatment assignment is independent of all potential variables. In stratified randomized experiments, there may be baseline covariates. For identification, the analysis should be conducted in each level of covariates. Assumption 2 rules out defiers. The treatment assignment (encouragement) should have non-negative effect on the actual treatment for each individual. Assumption 3 states that the treatment assignment can only exert its effect on survival or outcome through the actual treatment. There is no direct effect of the treatment assignment on the survival or outcome. Under Assumption 3, we can write $S(z,d)=S(d)$ and $Y(z,d)=Y(d)$. Assumption 4 ensures that compliers, who would always adhere to the treatment assignment, must exist.
Stratified by the potential values of $(D(0),D(1))$ and $(S(0),S(1))$, the whole population can be divided into eight principal strata, as listed in Table 1. There is only one principal stratum, the survived compliers ($cl$), in which a meaningful causal estimand can be defined. In other principal strata, at least one potential outcome $Y(z)$ ($z=0,1$) is undefined. To evaluate the treatment effect, we should restrict the target population to the survived compliers. Therefore, we define the principal average causal effect (PACE) in the survived compliers as
$\tau=E\{Y(1)-Y(0)\mid G=cl\},$
which measures the magnitude of average treatment effects in compliers who would always survive regardless of the treatment received.
Table 1: Principal strata in the population. The notation ✓is used to denote well-definedness and $\times$ to denote undefinedness
Stratum $D(1)$ $D(0)$ $S(1)$ $S(0)$ $Y(1)$ $Y(0)$ Description
$al$ 1 1 1 $\times$ $\times$ Survived always-takers
$ad$ 1 1 0 $\times$ $\times$ $\times$ Dead always-takers
$cl$ 1 0 1 1 Survived compliers
$cp$ 1 0 1 0 $\times$ Protected compliers
$ch$ 1 0 0 1 $\times$ Harmed compliers
$cd$ 1 0 0 0 $\times$ $\times$ Doomed compliers
$nl$ 0 0 $\times$ 1 $\times$ Survived never-takers
$nd$ 0 0 $\times$ 0 $\times$ $\times$ Dead never-takers
Some outcomes are truncated by death and missingness. We assume that $Y$ will be missing if $S$ is missing. To account for truncation-by-death (the problem of partially defined outcomes) and unintended missingness, the assumptions below are also required for identification.
Assumption 5 (Principal ignorability)
$E\{Y(1)\mid G=cl\}=E\{Y(1)\mid G=cp\}$ and $E\{Y(0)\mid G=cl\}=E\{Y(0)\mid G=ch\}$.
Assumption 6 (Missingness at random)
[…]
Assumption 5 means that the potential outcome of interest $Y(d)$ does not have cross-world reliance on $S(1-d)$ as long as $S(d)=1$ (Jo and Stuart, , 2009). This assumption is sometimes referred as explainable nonrandom survival (Hayden et al., , 2005; Egleston et al., , 2007, 2009) to indicate that the survival cannot causally predict the potential outcome of interest. For notation simplicity, denote the compliance strata
$a$ =$=\{D(1)=D(0)=1\}=\{al,ad\},$
$c$ =$=\{D(1)>D(0)\}=\{cl,cp,ch,cd\},$
$n$ =$=\{D(1)=D(0)=0\}=\{nl,nd\}.$
Let $p_{g}=P(G=g)$ be the proportion of principal stratum $g$. Assumption 5 merges always-survivors and protected subpopulations in the treated group ($cl$ and $cp$), and merges always-survivors and harmed populations in the control group ($cl$ and $ch$) for compliers. Under Assumption 5,
$E\{Y(1)\mid G=cl\}$ $=E\{Y(1)\mid G=cp\}=E\{Y(1)\mid S(1)=1,c\},$
$E\{Y(0)\mid G=cl\}$ $=E\{Y(0)\mid G=ch\}=E\{Y(0)\mid S(0)=1,c\},$
$E\{Y(1)\mid G=al\}$ $=E\{Y(1)\mid S(1)=1,a\},$
$E\{Y(0)\mid G=nl\}$ $=E\{Y(0)\mid S(0)=1,n\}.$
Assumption 6 means that the missingness of an variable is independent of potential outcomes given observed data prior to this variable. Although $Y(d)$ is not always well defined, the product $Y(d)S(d)$ can always be well defined. Under Assumptions 5, 6 and consistency, for $z,d=0,1$, we have
$E\{Y(d)S(d)\mid Z=z,D=d\}$ $=E\{S\mid Z=z,D=d,\Delta_{S}=1\}$
$\quad\cdot E\{Y\mid Z=z,D=d,S=1,\Delta_{Y}=1\}.$
So we can identify the mean product of the survival and outcome in any subgroup stratified by the assigned treatment and received treatment. This result allows us to estimate the conditional expectations of $S$ and $Y$ separately using available data, improving the efficiency of estimation.
2.2 Identification
The first-stage regression gives the proportions of compliance groups. The individuals receiving treatment come from always-takers and compliers in the treated group, and the individuals receiving treatment come solely from always-takers in the control group. Therefore, we have
$E\{D\mid Z=1\}$ $=p_{a}+p_{c},$
$E\{D\mid Z=0\}$ $=p_{a},$
from which we can solve
$p_{a}$ $=E\{D\mid Z=0\},$
$p_{c}$ $=E\{D\mid Z=1\}-E\{D\mid Z=0\},$
$p_{n}$ $=1-E\{D\mid Z=1\}.$
The second-stage regression gives the proportions of survival groups in compliers. Solving the following equations,
$E\{S\mid Z=1,D=1\}$ $=\frac{p_{a}}{p_{a}+p_{c}}P(S(1)=1\mid a)+\frac{p_{c}}{p_{a}+p_{c}}P(S(1)=1\mid c),$
$E\{S\mid Z=1,D=0\}$ $=P(S(0)=1\mid n),$
$E\{S\mid Z=0,D=1\}$ $=P(S(1)=1\mid a),$
$E\{S\mid Z=0,D=0\}$ $=\frac{p_{c}}{p_{c}+p_{n}}P(S(0)=1\mid c)+\frac{p_{n}}{p_{c}+p_{n}}P(S(0)=1\mid n),$
we obtain
$P(S(1)=1\mid c)$ =$=\frac{\{p_{c}+p_{a}\}E\{S\mid Z=1,D=1\}-p_{a}E\{S\mid Z=0,D=1\}}{p_{c}},$
$P(S(0)=1\mid c)$ =$=\frac{\{p_{c}+p_{n}\}E\{S\mid Z=0,D=0\}-p_{n}E\{S\mid Z=1,D=0\}}{p_{c}}$
The third-stage regression links the outomes with principal strata. Note that $YS$ is always well defined, so we use $YS$ as the outcome.
$E\{YS\mid Z=1,D=1\}$ $=\frac{p_{a}}{p_{a}+p_{c}}P(S(1)=1\mid a)E\{Y(1)\mid S(1)=1,a\}$
$\quad+\frac{p_{c}}{p_{a}+p_{c}}P(S(1)=1\mid c)E\{Y(1)\mid S(1)=1,c\},$
$E\{YS\mid Z=1,D=0\}$ $=P(S(0)=1\mid n)E\{Y(0)\mid S(0)=1,n\},$
$E\{YS\mid Z=0,D=1\}$ $=P(S(1)=1\mid a)E\{Y(1)\mid S(1)=1,a\},$
$E\{YS\mid Z=0,D=0\}$ $=\frac{p_{c}}{p_{c}+p_{n}}P(S(0)=1\mid c)E\{Y(0)\mid S(0)=1,c\}$
$\quad+\frac{p_{n}}{p_{c}+p_{n}}P(S(0)=1\mid n)E\{Y(0)\mid S(0)=1,n\}.$
The solution of the above equations is
$E\{Y(1)\mid G=cl\}$ =$=\frac{\{p_{a}+p_{c}\}E\{YS\mid Z=1,D=1\}-p_{a}E\{YS\mid Z=0,D=1\}}{p_{c}P(S(1)=1\mid c)}$
=$=\frac{\{p_{a}+p_{c}\}E\{YS\mid Z=1,D=1\}-p_{a}E\{YS\mid Z=0,D=1\}}{\{p_{a}+p_{c}\}E\{S\mid Z=1,D=1\}-p_{a}E\{S\mid Z=0,D=1\}}$
[…]
$E\{Y(0)\mid G=cl\}$ =$=\frac{\{p_{c}+p_{n}\}E\{YS\mid Z=0,D=0\}-p_{n}E\{YS\mid Z=1,D=0\}}{p_{c}P(S(0)=1\mid c)}$
=$=\frac{\{p_{c}+p_{n}\}E\{YS\mid Z=0,D=0\}-p_{n}E\{YS\mid Z=1,D=0\}}{\{p_{c}+p_{n}\}E\{S\mid Z=0,D=0\}-p_{n}E\{S\mid Z=1,D=0\}}$
[…]
$E\{Y(1)\mid G=al\}$ $=\frac{E\{YS\mid Z=0,D=1\}}{P(S(1)=1\mid a)}$
$=\frac{E\{S\mid(1-Z)D\Delta_{S}=1\}E\{Y\mid(1-Z)DS\Delta_{Y}=1\}}{E\{S\mid Z=0,D=1\}},$
$E\{Y(0)\mid G=nl\}$ $=\frac{E\{YS\mid Z=1,D=0\}}{P(S(0)=1\mid n)}$
$=\frac{E\{S\mid Z(1-D)\Delta_{S}=1\}E\{Y\mid Z(1-D)S\Delta_{Y}=1\}}{E\{S\mid Z=1,D=0\}}.$
Afterwards, the identifiability result follows immediately.
Theorem 1
Under Assumptions 16, the expectation of potential outcomes in survived compliers $E\{Y(d)\mid cl\}$ and the principal average causal effect $\tau$ are nonparametrically identifiable.
When the way of breaking randomized experiments is simple, we can simplify the identification results.
Remark 1
If there is no truncation-by-death, i.e., $P(S(0)=1\mid D(1),D(0))=1$, then the principal average causal effect is reduced to the Wald estimator in instrumental variables,
$\tau$ $=\frac{E\{Y\mid Z=1,\Delta_{Y}=1\}-E\{Y\mid Z=0,\Delta_{Y}=1\}}{E\{D\mid Z=1\}-E\{D\mid Z=0\}}.$
Remark 2
If there is no non-compliance, i.e., $P(D(1)>D(0))=1$, then the principal average causal effect is reduced to the observed survivor-case contrast,
$\tau=E\{Y\mid Z=1,S=1,\Delta_{Y}=1\}-E\{Y\mid Z=0,S=1,\Delta_{Y}=1\}.$
In some circumstances, the principal ignorability (Assumption 5) may seem strong. However, the principal average causal effect would not be identifiable without introducing additional information on the cross-world reliance of $Y$ on $S$ or invoking a substitutional variable for the principal stratum $G$ if the principal ignorability assumption is violated.
Remark 3
If there is no missingness, then the principal average causal effect is
$\tau$ $=\frac{E\{DSY\mid Z=1\}-E\{DSY\mid Z=0\}}{E\{DS\mid Z=1\}-E\{DS\mid Z=0\}}$
$\quad-\frac{E\{(1-D)SY\mid Z=1\}-E\{(1-D)SY\mid Z=0\}}{E\{(1-D)S\mid Z=1\}-E\{(1-D)S\mid Z=0\}}.$
In the presence of missingness, we can simply use imputation methods to impute the missing values, and use the formula in Remark 3 to estimate the principal average causal effect (Harel and Zhou, , 2007; Kenward and Carpenter, , 2007; Rubin, , 2018). To be noted, the proportions of principal strata $cl$, $cp$, $ch$, $cd$ defined by compliance type and survival type are not identifiable without additional assumptions. A typical assumption is the individual-level monotonicity on $S$, i.e., $S(1)\geq S(0)$ almost surely, under which
$P(G=cl)=p_{c}P(S(0)=1\mid c)=E\{(1-D)S\mid Z=0\}-E\{(1-D)S\mid Z=1\}.$
Whether this assumption is plausible depends on specific scientific questions. For example, this monotonicity does not hold for JC data (Frumento et al., , 2012) but may be reasonable in some medical applications (Deng et al., , 2021).
The unidentifiability of principal strata proportions brings difficulty in aggregating the conditional causal effect in each covariates level to the average causal effect in the overall population. In the presence of baseline covariates $X$, we can impose the identification assumptions in each level of $X$ and identify the conditional principal average causal effect (CPACE)
$\tau(x)=E\{Y(1)-Y(0)\mid G=cl,X=x\}.$
However, since we do not know $P(G=cl\mid X=x)$, we can not aggregate the condition effects
[…]
Therefore, the principal stratification analysis on broken randomized experiments should be conducted in each level of covariates. Otherwise we need to impose additional assumptions; see Appendix for more details.
3 Estimation and inference
Consider a completely randomized design with known treatment propensity $P(Z=1)=1-P(Z=0)=r$. Suppose there is no missingness, otherwise one can use multiple imputation or other methods to generate a completed dataset first. The asymptotic variance of the estimator can be calculated by Rubin’s rule from completed datasets (Rubin and Schenker, , 1986; White et al., , 2011; Rubin, , 2018).
Let $n_{zds}=\sum_{i=1}^{n}1(Z_{i}=z,D_{i}=d,S_{i}=s)$ for $z,d,s=0,1$, and use dot $\cdot$” to represent summation with respect to its/their corresponding argument(s). Let $\mu_{d}=E\{Y(d)\mid G=cl\}$ for $d=0,1$. Denote
$\theta_{1}=E\{D\mid Z=1\},\ \theta_{11}=E\{S\mid ZD=1\},\ \theta_{10}=E\{S\mid Z(1-D)=1\},$
$\theta_{0}=E\{D\mid Z=0\},\ \theta_{01}=E\{S\mid(1-Z)D=1\},\ \theta_{00}=E\{S\mid(1-Z)(1-D)=1\},$
and let $f(y;\theta_{zd1})$ be the density function of $Y$ given $Z=z$, $D=d$ and $S=1$ with mean $b(\theta_{zd1})=E\{Y\mid Z=z,D=d,S=1\}$. Suppose there are no common parameters in $\theta_{zd1}$ for $z,d=0,1$. Write the unknown parameter as
$\theta=(\theta_{1},\theta_{0},\theta_{11},\theta_{10},\theta_{01},\theta_{00},\theta_{111},\theta_{101},\theta_{011},\theta_{001})^{{\mathrm{\scriptscriptstyle T}}}.$
The likelihood with respect to $\theta$ is
$L(\theta;\mathcal{O})$ $=(\theta_{1})^{n_{11\cdot}}(1-\theta_{1})^{n_{10\cdot}}\cdot(\theta_{11})^{n_{111}}(1-\theta_{11})^{n_{110}}\cdot(\theta_{10})^{n_{101}}(1-\theta_{10})^{n_{100}}$
$\quad\cdot(\theta_{0})^{n_{01\cdot}}(1-\theta_{0})^{n_{00\cdot}}\cdot(\theta_{01})^{n_{011}}(1-\theta_{01})^{n_{010}}\cdot(\theta_{00})^{n_{001}}(1-\theta_{00})^{n_{000}}$
$\quad\cdot\prod_{i:S_{i}=1}f(y_{i};\theta_{Z_{i}D_{i}1})\cdot r^{n_{1\cdot\cdot}}(1-r)^{n_{0\cdot\cdot}}.$
Solving the maximum likelihood estimation equation $\partial\log L(\theta;\mathcal{O})/\partial\theta=0$, one can obtain
$\widehat{\theta}=\left(\frac{n_{11\cdot}}{n_{1\cdot\cdot}},\frac{n_{01\cdot}}{n_{0\cdot\cdot}},\frac{n_{111}}{n_{11\cdot}},\frac{n_{101}}{n_{10\cdot}},\frac{n_{011}}{n_{01\cdot}},\frac{n_{001}}{n_{00\cdot}},\widehat{\theta}_{111},\widehat{\theta}_{101},\widehat{\theta}_{011},\widehat{\theta}_{001}\right)^{{\mathrm{\scriptscriptstyle T}}},$
where $\widehat{\theta}_{zd1}$ solves $\sum_{Z_{i}=z,D_{i}=d,S_{i}=1}\partial\log f(y_{i};\theta_{zd1})/\partial\theta_{zd1}=0$, with
$\operatorname{cov}(\widehat{\theta})$ $=\operatorname{diag}\bigg{\{}\frac{r}{n\theta_{1}(1-\theta_{1})},\frac{1-r}{n\theta_{0}(1-\theta_{0})},\frac{r\theta_{1}}{n\theta_{11}(1-\theta_{11})},\frac{r(1-\theta_{1})}{n\theta_{10}(1-\theta_{10})},$
$\qquad\qquad\frac{(1-r)\theta_{0}}{n\theta_{01}(1-\theta_{01})},\frac{(1-r)(1-\theta_{0})}{n\theta_{00}(1-\theta_{00})},\operatorname{var}(\widehat{\theta}_{111}),\operatorname{var}(\widehat{\theta}_{101}),\operatorname{var}(\widehat{\theta}_{011}),\operatorname{var}(\widehat{\theta}_{001})\bigg{\}}.$
If we use the sample average $\widehat{\theta}_{zd1}$ of the outcomes in $\{Z=z,D=d,S=1\}$ to estimate $b(\theta_{zd1})$, then the variance $\operatorname{var}(\widehat{\theta}_{zd1})$ is the standard deviation of the outcomes in $\{Z=z,D=d,S=1\}$ divided by the group size $n_{zd1}$.
Since we can express
$\mu_{1}=\frac{\theta_{1}\theta_{11}b(\theta_{111})-\theta_{0}\theta_{01}b(\theta_{011})}{\theta_{1}\theta_{11}-\theta_{0}\theta_{01}},\ \mu_{0}=\frac{(1-\theta_{1})\theta_{10}b(\theta_{101})-(1-\theta_{0})\theta_{00}b(\theta_{001})}{(1-\theta_{1})\theta_{10}-(1-\theta_{0})\theta_{00}},$
the causal paratemers can be estimated by plugging $\widehat{\theta}$ into $\mu_{d}$, denoted by $\widehat{\mu}_{d}$, for $d=0,1$. Similarly, the principal average causal effect is estimated by $\widehat{\tau}=\widehat{\mu}_{1}-\widehat{\mu}_{0}$. The asymptotic properties of the plug-in estimators can be established by sandwich estimator, as shown in the following theorem.
Theorem 2
Under Assumptions 15, suppose there is no missingness. Let $\Gamma_{d}(\theta)=\partial\mu_{d}/\partial\theta$ for $d=0,1$, then
$n^{1/2}\left\{\widehat{\mu}_{1}-\mu_{1}\right\}\rightarrow_{d}N\left\{0,\ n\Gamma_{1}(\theta)^{{\mathrm{\scriptscriptstyle T}}}\operatorname{cov}(\widehat{\theta})\Gamma_{1}(\theta)\right\},$
$n^{1/2}\left\{\widehat{\mu}_{0}-\mu_{0}\right\}\rightarrow_{d}N\left\{0,\ n\Gamma_{0}(\theta)^{{\mathrm{\scriptscriptstyle T}}}\operatorname{cov}(\widehat{\theta})\Gamma_{0}(\theta)\right\},$
$n^{1/2}\left\{\widehat{\tau}-\tau\right\}\rightarrow_{d}N\left\{0,\ n\left(\Gamma_{1}(\theta)-\Gamma_{0}(\theta)\right)^{{\mathrm{\scriptscriptstyle T}}}\operatorname{cov}(\widehat{\theta})\left(\Gamma_{1}(\theta)-\Gamma_{0}(\theta)\right)\right\}.$
When the outcome $Y$ is binary, the normal approximation may not be good, since the distribution of the estimator $\widehat{\mu}_{d}$ can be skewed. To respect the fact that $\mu_{d}$ must lie in $[0,1]$, we may as well adopt $\operatorname{logit}(\mu_{d})=\log\{\mu_{d}/(1-\mu_{d})\}$ as the target estimand, which can range on the whole real line. The contrast $\tau^{*}=\operatorname{logit}(\mu_{1})-\operatorname{logit}(\mu_{0})$ reflects the log odds ratio between treatment and control for survived compliers. However, the log odds ratio does not enjoy collapsibility. Since $Y$ is binary, we have $\theta_{zd1}=E\{Y\mid Z=z,D=d,S=1\}$. Now
$\operatorname{logit}(\mu_{1})$ $=\log\left\{\frac{\theta_{1}\theta_{11}\theta_{111}-\theta_{0}\theta_{01}\theta_{011}}{\theta_{1}\theta_{11}(1-\theta_{111})-\theta_{0}\theta_{01}(1-\theta_{011})}\right\},$
$\operatorname{logit}(\mu_{0})$ $=\log\left\{\frac{(1-\theta_{1})\theta_{10}\theta_{101}-(1-\theta_{0})\theta_{00}\theta_{001}}{(1-\theta_{1})\theta_{10}(1-\theta_{101})-(1-\theta_{0})\theta_{00}(1-\theta_{001})}\right\}.$
Plugging $\widehat{\theta}$ into the expression of $\tau^{*}$, denoted by $\widehat{\tau}^{*}$, it is straightforward to obtain a similar result with Theorem 2. We can use the delta method to obtain the asymptotic variance for the estimator of $\operatorname{logit}(\mu_{d})$ and construct the confidence interval. The we can use the delta method again to obtain the asymptotic variance for the estimator of $\mu_{d}$ and the principal average causal effect $\tau$. Using the method of variance estimates recovery or delta method (Donner and Zou, , 2011), we can get an asymmetric confidence interval for $\tau$ which usually has better coverage rates in finite sample.
There is another equivalent approach to establish the asymptotic property for the logit transformed mean outcomes. Let $p_{zdy}=P(Z=z,D=d,S=1,Y=y)$, $\widehat{p}_{zdy}=n^{-1}\sum_{i=1}^{n}1(Z_{i}=z,D_{i}=d,S_{i}=1,Y_{i}=y)$ for $z,d,y=0,1$, and use dot $\cdot$” to represent summation with respect to its/their corresponding argument(s).
Under the identification assumptions, the point estimates of the average potential outcomes in the survived compliers are given by
$\widehat{\mu}_{1}=\frac{\frac{\widehat{p}_{111}}{r}-\frac{\widehat{p}_{011}}{1-r}}{\frac{\widehat{p}_{11\cdot}}{r}-\frac{\widehat{p}_{01\cdot}}{1-r}},\quad\widehat{\mu}_{0}=\frac{\frac{\widehat{p}_{101}}{r}-\frac{\widehat{p}_{001}}{1-r}}{\frac{\widehat{p}_{10\cdot}}{r}-\frac{\widehat{p}_{00\cdot}}{1-r}},$
or by logit transformation,
$\operatorname{logit}(\widehat{\mu}_{1})$ $=\log\left(\frac{\widehat{p}_{111}}{r}-\frac{\widehat{p}_{011}}{1-r}\right)-\log\left(\frac{\widehat{p}_{110}}{r}-\frac{\widehat{p}_{010}}{1-r}\right),$
$\operatorname{logit}(\widehat{\mu}_{0})$ $=\log\left(\frac{\widehat{p}_{101}}{r}-\frac{\widehat{p}_{001}}{1-r}\right)-\log\left(\frac{\widehat{p}_{100}}{r}-\frac{\widehat{p}_{000}}{1-r}\right).$
Theorem 3
Under Assumptions 15, suppose there is no missingness. Then
$n^{1/2}\left\{\operatorname{logit}(\widehat{\mu}_{1})-\operatorname{logit}(\mu_{1})\right\}\rightarrow_{d}N(0,V_{1}),\quad n^{1/2}\left\{\operatorname{logit}(\widehat{\mu}_{0})-\operatorname{logit}(\mu_{0})\right\}\rightarrow_{d}N(0,V_{0}),$
$n^{1/2}\left\{\tau^{*}-\tau\right\}\rightarrow_{d}N(0,V_{1}+V_{0}-2V_{10}),$
$n^{1/2}\left\{\widehat{\mu}_{1}-\mu_{1}\right\}\rightarrow_{d}N(0,\mu_{1}^{2}(1-\mu_{1})^{2}V_{1}),\quad n^{1/2}\left\{\widehat{\mu}_{0}-\mu_{0}\right\}\rightarrow_{d}N(0,\mu_{0}^{2}(1-\mu_{0})^{2}V_{0}),$
$n^{1/2}\left\{\widehat{\tau}_{0}-\tau_{0}\right\}\rightarrow_{d}N(0,\mu_{1}^{2}(1-\mu_{1})^{2}V_{1}+\mu_{0}^{2}(1-\mu_{0})^{2}V_{0}-2\mu_{1}\mu_{0}(1-\mu_{1})(1-\mu_{0})V_{10}),$
where
$V_{1}$ $=\frac{\frac{p_{111}(1-2p_{111})}{r^{2}}+\frac{p_{011}}{(1-r)^{2}}}{(\frac{p_{111}}{r}-\frac{p_{011}}{1-r})^{2}}+\frac{\frac{p_{110}(1-2p_{110})}{r^{2}}+\frac{p_{010}}{(1-r)^{2}}}{(\frac{p_{110}}{r}-\frac{p_{010}}{1-r})^{2}}+\frac{\frac{4p_{111}p_{110}}{r^{2}}}{(\frac{p_{111}}{r}-\frac{p_{011}}{1-r})(\frac{p_{110}}{r}-\frac{p_{010}}{1-r})},$
$V_{0}$ $=\frac{\frac{p_{101}(1-2p_{101})}{r^{2}}+\frac{p_{001}}{(1-r)^{2}}}{(\frac{p_{101}}{r}-\frac{p_{001}}{1-r})^{2}}+\frac{\frac{p_{100}(1-2p_{100})}{r^{2}}+\frac{p_{000}}{(1-r)^{2}}}{(\frac{p_{100}}{r}-\frac{p_{000}}{1-r})^{2}}+\frac{\frac{4p_{101}p_{110}}{r^{2}}}{(\frac{p_{101}}{r}-\frac{p_{001}}{1-r})(\frac{p_{100}}{r}-\frac{p_{000}}{1-r})},$
$V_{10}$ $=-\frac{\frac{2(p_{101}p_{000}-p_{100}p_{001})(p_{111}p_{010}-p_{110}p_{011})}{r^{2}(1-r)^{2}}}{(\frac{p_{111}}{r}-\frac{p_{011}}{1-r})(\frac{p_{110}}{r}-\frac{p_{010}}{1-r})(\frac{p_{101}}{r}-\frac{p_{001}}{1-r})(\frac{p_{100}}{r}-\frac{p_{000}}{1-r})}.$
By logit transformation, the $(1-\alpha)100\%$ asymmetric confidence intervals for $\mu_{1}$ and $\mu_{0}$ can be constructed as
$(L_{1},U_{1})$ $=\operatorname{expit}\left\{\operatorname{logit}(\widehat{\mu}_{1})\pm z_{1-\alpha/2}(V_{1}/n)^{1/2}\right\},$
$(L_{0},U_{0})$ $=\operatorname{expit}\left\{\operatorname{logit}(\widehat{\mu}_{0})\pm z_{1-\alpha/2}(V_{0}/n)^{1/2}\right\}$
respectively, where $\operatorname{expit}(x)=\exp(x)/\{1+\exp(x)\}$ and $z_{1-\alpha/2}$ is the upper $\alpha/2$ quantile of standard normal distribution. The $(1-\alpha)100\%$ asymmetric confidence interval for $\tau$ can be constructed as $(L,U)$, where
$L=\widehat{\tau}-\left\{(L_{1}-\widehat{\mu}_{1})^{2}-(U_{0}-\widehat{\mu}_{0})^{2}-2\widehat{\mu}_{0}\widehat{\mu}_{1}(1-\widehat{\mu}_{1})(1-\widehat{\mu}_{0})(L_{1}-\widehat{\mu}_{1})(U_{0}-\widehat{\mu}_{0})\right\}^{1/2},$
$U=\widehat{\tau}+\left\{(U_{1}-\widehat{\mu}_{1})^{2}-(L_{0}-\widehat{\mu}_{0})^{2}-2\widehat{\mu}_{0}\widehat{\mu}_{1}(1-\widehat{\mu}_{1})(1-\widehat{\mu}_{0})(U_{1}-\widehat{\mu}_{1})(L_{0}-\widehat{\mu}_{0})\right\}^{1/2}.$
For continuous $Y$, one can define a new outcome $Y_{c}=1_{Y>c}$ given any arbitrary threshold $c\in\mathbb{R}$. In this manner, one can obtain the estimators for the cumulative distribution functions of $Y(1)$ and $Y(0)$ in the $cl$ stratum with pointwise $(1-\alpha)100\%$ confidence intervals by varying the threshold $c$ from $\infty$- ∞ to $\infty$.
Although it is possible to derive the asymptotic properties with the existence of missingness, the expression of the asymptotic variance would be too complicated and it would be inconvenient for practical use. Therefore, the discussion with missing data is omitted here. We recommend practitioners adopt off-the-shelf tools like multiple imputation to disentangle missing data. Analyses can be performed for each completed dataset, and inference is trivial to combine these results using Rubin’s rule (Rubin, , 2018).
4 Simulation
In this section, we conduct a simulation study to assess the performance of the proposed method. Suppose that there are $n\in\{500,2000,8000\}$ independent individuals in the sample. Each individual is assigned to the treated group or control group with equal probability $r=0.5$. The compliance type is generated as
$P(D(0)=1)=0.3,\ P(D(1)=1\mid D(0)=0)=0.4.$
So the proportion of always-takers, compliers and never-takers is 0.3, 0.28 and 0.42, respectively. The survival type is generated as
$P(S(0)=1\mid D(0),D(1))$ $=0.3+0.2D(0)+0.2D(1),$
$P(S(1)=1\mid D(0),D(1))$ $=0.3+0.3D(0)+0.3D(1).$
To generate the outcome, we first assume that the distribution of potential outcomes are homogeneous among compliance types (Case 1). That is, $Y(d)$ does not depend on $(D(1),D(0))$. As long as $S(d)=1$, we generate
$\{Y(d)\mid S(d)=1\}\sim N(1+d,(0.8+0.2d)^{2}).$
Next, we consider another scenario that the distribution of potential outcomes are heterogeneous among compliance types (Case 2). For compliers,
$\{Y(d)\mid G=c,S(d)=1\}\sim N(1+d,(0.8+0.2d)^{2}).$
For never-takers and always-takers,
$\{Y(0)\mid G=n,S(0)=1\}$ $\sim N(1,0.8^{2})-N(0.5,0.2^{2}),$
$\{Y(1)\mid G=a,S(1)=1\}$ $\sim N(2,1.0^{2})+N(0.3,0.2^{2}).$
The observed variables under consistency
$D=ZD(1)+(1-Z)D(0),S=DS(1)+(1-D)S(0),Y=DY(1)+(1-D)Y(0).$
We assume there is no missingness. Assumptions 15 hold under the data generating process.
To assess the robustness of the proposed method against principal ignorability (Assumption 5), which is the assumption most likely to be violated in practice, we further consider two scenarios. We assume that the potential outcome $Y(d)$ has cross-world reliance on $S(1-d)$. In Case 3, we add $S(1)$ to $Y(0)$ and add $0.5S(0)$ to $Y(1)$ under homogeneity (Case 1). In Case 4, we add $S(1)$ to $Y(0)$ and add $0.5S(0)$ to $Y(1)$ under heterogeneity (Case 2). The true treatment effect is empirically evaluated by the mean difference in $Y(1)$ and $Y(0)$ among survived compliers $\{D(1)>D(0),S(1)=S(0)=1\}$.
We compare two estimators. The first estimator is the proposed one (denoted by PACE), and the standard error of the treatment effect estimator can be calculated by Theorem 2. The second estimator is the two-stage least square estimator (denoted by 2SLS), where we restrict the second-stage regression in the sample of observed survivors $\{S=1\}$. Table 2 shows the bias, standard deviation (SD), standard error (SE) and coverage rate of the nominal 95% confidence interval. The proposed method PACE has ignorable bias in Cases 1 and 2 where all assumptions hold. When principal ignorability is violated, the proposed method is biased. The 2SLS estimator is biased in all settings because it cannot deal with the cross-world reliance of $S(d)$ on $(D(1),D(0))$. With the sample size being larger, the standard deviation and standard error get smaller.
Table 2: The bias, standard deviation (SD), standard error (SE) and coverage rate of the nominal 95% confidence interval.
Principal ignorability Not principal ignorability
Homogeneous Heterogeneous Homogeneous Heterogeneous
Size Method PACE 2SLS PACE 2SLS PACE 2SLS PACE 2SLS
500 Bias 0.004 0.188 -0.069 0.115 0.141 0.296 0.068 0.223
SD 0.393 0.490 0.504 0.621 0.433 0.505 0.568 0.642
SE 0.425 0.502 0.608 0.658 0.475 0.521 0.693 0.666
Coverage 0.982 0.945 0.972 0.958 0.960 0.925 0.941 0.940
2000 Bias 0.003 0.183 -0.012 0.136 0.148 0.285 0.133 0.238
SD 0.179 0.230 0.216 0.293 0.198 0.240 0.242 0.303
SE 0.184 0.244 0.255 0.319 0.203 0.253 0.288 0.323
Coverage 0.964 0.895 0.984 0.944 0.884 0.796 0.924 0.893
8000 Bias -0.003 0.180 -0.005 0.142 0.146 0.283 0.144 0.245
SD 0.090 0.118 0.111 0.151 0.101 0.125 0.126 0.160
SE 0.089 0.122 0.124 0.159 0.099 0.126 0.140 0.161
Coverage 0.943 0.701 0.968 0.858 0.670 0.379 0.810 0.659
5 Application to Job Corps data
The U.S. Job Corps (JC) training program was a randomized education and job training program for disadvantaged youths. The study collected baseline covariates including gender, age, race, education status and family information. There are 9240 complete cases, among which 5577 were assigned to Job Corps and 3663 out of Job Corps. In the first year after assignment, 84.6% enrolled in education and/or vocational training in the treated group, and 50.6% enrolled in education and/or vocational training in the control group. In the second year after assignment, 52.3% enrolled in education and/or vocational training in the treated group, and 42.0% enrolled in education and/or vocational training in the control group. In our analysis, we regard the individual as receiving the training if he/she has ever enrolled in the training when analyzing the treatment effects later than the second year after assignment. So 88.8% in the treated group actually received the treatment and 60.6% in the control group actually received the treatment. We infer that the proportion of always-takers, compliers and never-takers is 0.606, 0.282 and 0.112, respectively.
In one year after assignment, 49.4% have ever worked in the treated group, and 57.1% have ever worked in the control group. In the second year after assignment, the proportion of employment is 75.2% in the treated group and 75.3% in the control group. In the third year after assignment, the proportion of employment increases to 83.5% in the treated group and 81.4% in the control group. In the fourth year after assignment, the proportion of employment increases to 83.7% in the treated group and 81.3% in the control group. We see that the employment status becomes stable in three years after assignment. We take the logarithm for the weekly earnings. There is a trend that the earnings grow higher with time going on. Table 3 lists the proportion of employment and average earnings in each level of the assigned treatment $Z$ and received treatment $D$.
Table 3: The proportion of employment and average earnings in each level of the assigned treatment $Z$ and received treatment $D$ in the Job Corps study
Year Employment $S$ Earnings $Y$
$(Z,D)$ (1,1) (1,0) (0,1) (0,0) (1,1) (1,0) (0,1) (0,0)
1 0.480 0.574 0.543 0.601 4.824 5.090 4.803 5.041
2 0.749 0.772 0.745 0.766 4.723 4.841 4.661 4.868
3 0.838 0.812 0.807 0.824 5.023 4.964 4.924 5.009
4 0.840 0.815 0.827 0.792 5.207 5.128 5.160 5.168
Figure 1 shows the estimated treatment effect on employment $E\{S(1)-S(0)\mid c\}$ and the estimated treatment effect on weekly earnings $E\{Y(1)-Y(0)\mid cl\}$ with 95% confidence intervals. In the first year after assignment, there is a negative effect on employment. This is possibly because the individuals spent their time on training and they were not ready to enter the job market in the first year. Since the third year after assignment, there is a positive effect on employment, indicating that the training program helped the youths finding jobs. However, the effect is not significant at level 0.05. For the always-employed people, the training program has a significant effect on earnings.
Figure 1: Time-varying treatment effect on employment $E\{S(1)-S(0)\mid c\}$ and time-varying treatment effect on earnings $E\{Y(1)-Y(0)\mid cl\}$.
Table 4: The estimated mean outcomes, treatment effect, standard error (se) and $P$-value in different years by the proposed method (PACE), two-stage least square (2SLS), intention-to-treat analysis (ITT), as-treated analysis (AT) and per-protocol analysis (PP) for survived compliers
Year Method $\widehat{\mu}_{1}$ se $\widehat{\mu}_{0}$ se $\widehat{\tau}$ se $P$
1 PACE 4.902 (0.090) 5.008 (0.043) -0.106 (0.101) 0.291
2SLS -0.128 (0.080) 0.107
ITT 4.859 (0.018) 4.903 (0.020) -0.044 (0.027) 0.105
AT 4.800 (0.016) 5.040 (0.023) -0.240 (0.028) 0.000
PP 4.813 (0.020) 5.027 (0.027) -0.214 (0.034) 0.000
2 PACE 4.855 (0.082) 4.879 (0.050) -0.025 (0.096) 0.798
2SLS -0.026 (0.095) 0.788
ITT 4.737 (0.017) 4.744 (0.021) -0.007 (0.027) 0.789
AT 4.704 (0.015) 4.860 (0.027) -0.156 (0.031) 0.000
PP 4.723 (0.018) 4.868 (0.033) -0.145 (0.037) 0.000
3 PACE 5.211 (0.068) 5.027 (0.046) 0.185 (0.081) 0.022
2SLS 0.207 (0.082) 0.012
ITT 5.016 (0.014) 4.958 (0.019) 0.058 (0.023) 0.013
AT 4.993 (0.013) 4.996 (0.025) -0.003 (0.028) 0.921
PP 5.023 (0.015) 5.009 (0.030) 0.013 (0.034) 0.687
4 PACE 5.305 (0.067) 5.185 (0.047) 0.120 (0.081) 0.137
2SLS 0.127 (0.082) 0.121
ITT 5.199 (0.014) 5.163 (0.018) 0.036 (0.023) 0.121
AT 5.193 (0.013) 5.156 (0.025) 0.037 (0.028) 0.184
PP 5.207 (0.015) 5.168 (0.030) 0.039 (0.033) 0.239
Table 4 summarizes the estimated mean outcomes, treatment effect on weekly earnings, standard error (se) and $P$-value in different years by the proposed method (PACE), two-stage least square (2SLS), intention-to-treat analysis (ITT), as-treated analysis (AT) and per-protocol analysis (PP) for survived compliers. According to the proposed estimator, in the third year after assignment, there is a significant and positive effect on weekly earnings. The always-employed and compliers can have 18.5% higher salaries if taking the training program. The effect slightly goes smaller in the fourth year because the weekly earnings without participating in the training program $\mu_{0}$ got higher. Different methods give almost the same signs of the treatment effect. In the third and forth year, ITT, AT and PP underestimate the principal stratum treatment effect, indicating that compliers can benefit more from the training program. The proposed method yields the largest standard error among all the methods under comparison, because we define the estimand in a specific principal stratum, which has the fewest individuals among all methods. The point estimate of 2SLS is similar to that of PACE, indicating that the cross-world reliance of employment on the compliance type is weak (which is consistent with the left panel of Figure 1). The 2SLS has slightly smaller variance than PACE.
We can further incorporate covariates in the estimation. We choose six baseline covariates: gender, age, log household size, years of education, ever worked and health. There are missing values in covariates. We perform 10 times of imputation. In each time, we use simple imuptation to fill the missinge values by drawing values from the observed covariates distribution. Then we use linear regression to model the earnings $Y$ on covariates. Finally, we use the Rubin’s rule to calculate the sampling variance. Figure 2 shows the stimated treatment effect on earnings. Compared with Figure 1, we obtain tighter confidence intervals. The treatment effects are significant in the third and fourth years. The $P$-values are 0.0152, 0.6100, 0.0000 and 0.0036 for these four years, respectively. It is interesting that the treatment effect is significantly negative in the first year, indicating that joining the training program leads to lower earnings, probably because that the individuals in the program have shorter work time and hence less earnings. Figure 3 presents the estimated treatment effects by other methods with covariates adjustment. The trends are similar but the magnitudes differ. The confidence intervals given by 2SLS, ITT, AT and PP are tighter because these four methods do not account for the uncertainty in estimating the probability of survived (always-employed) compliers.
Figure 2: Time-varying treatment effect on earnings $E\{Y(1)-Y(0)\mid cl\}$, with covariates adjustment.
Figure 3: Time-varying treatment effect on earnings $E\{Y(1)-Y(0)\mid cl\}$ estimated by 2SLS, ITT, AT and PP, with covariates adjustment.
The results by the proposed method have meaning causal interpretations. However, since the target population, consisting of the survived compliers, is not individually identifiable, how to guide the results for future policy making is doubtful. Although we know the proportion of survived compliers, we do not know who will benefit from the training program. From the perspective of policy makers, the intention-to-treat analysis is the easiest to interpret, as it basically evaluates the effect of the treatment policy, regardless of whether the individuals adhere to the treatment assignment. In contrast, the proposed method evaluates a pure effect of the training program. The estimate by the proposed method is only valid for survived compliers, so we can never know how things would be if the training is made compulsory. Nevertheless, different methods approaches the scientific question from different perspectives, and assessment by different methods provides a comprehensive knowledge of how the treatment can make an effect.
6 Discussion
Post-treatment events bring challenges to conduct causal analysis of the treatment effect. Although the baseline characteristics are naturally balanced by randomization, these characteristics are not balanced anymore if there is post-treatment selection. In this paper, we consider the broken randomized experiments with non-compliance, truncation-by-death and missing data. Non-compliance leads the received treatment different from the assigned treatment, and the outcomes after the received treatment are still meaningful, reflecting a mixed effect of the treatment and possibly unmeasured compliance type. Truncation-by-death renders the outcomes ill defined or undefined. The survival type also embeds information about underlying characteristics of individuals. When there are non-compliance, truncation-by-death and missingness simultaneously in a randomized trial, we propose a framework to identify and estimate the principal stratum average treatment effect. The causal effect is defined in a subpopulation called the survived compliers. Although we can neither identify who belongs to this subpopulation or identify the proportion of this subpopulation, we are able to identify the treatment effect in this subpopulation.
There are some limitations for the estimand. First, the identification involves untestable assumptions. The first four assumptions on non-compliance are in line with the assumptions to identify the complier average causal effect (CACE) or the local average treatment effect (LATE). Lots of efforts have been devoted to find testable implications for these assumptions (Yang et al., , 2014; Mogstad et al., , 2018; Kédagni and Mourifié, , 2020). The principal ignorability assumption is hard to justify, and sensitivity analysis may be helpful to assess the robustness of substantial results to violation of principal ignorability. Utilizing additional information such as substitutional variables or post-treatment correlates, principal ignorability can be relaxed (Tchetgen Tchetgen, , 2014; Wang et al., , 2017).
Second, we did not consider adjustments of measured covariates, especially in the identification assumptions. More efficient estimator can be proposed by incorporating covariates in the modeling. However, relaxing ignorability assumptions by conditioning on covariates could incur other challenges because the average potential outcomes need to be integrated with respect to the distribution of covariates in the target principal stratum while the proportion of this principal stratum given covariates has not been identified. In Appendix, we provide some identification results with baseline covariates. This issue is important if we aim to extend the proposed framework to observational studies.
Third, as a common criticism to principal stratification, it is hard to generalize the result to a new population. Even if the policy is effective, policy makers cannot easily implement such a policy because we never know who will follow the policy instructions and benefit from the policy. As a tool, principal stratification defines the estimand in a subpopulation where all potential outcomes are well defined and comparabe. As a goal, the principal stratum shuold be consistent with the scientific interest. In order that the estimated treatment effect has a meaningful interpretation, the application scenario should be careful justified.
Data availability statement
The data we use in our work is publicly available in the R package causalweight. The R codes that support our findings are available as part of the online supplementary material.
Acknowledgements
We thank the support from the Save 2050 Programme jointly sponsored by Swarma Club and X-Order. JEL: C18, C39, C51
Funding information
This work is supported by the National Key Research and Development Program of China, Grant No. 2021YFF0901400; National Social Science Fund of China, Grant No. 19BJY108.
Appendix: Identification of the causal effect with covariates
For simplicity, suppose there are no missing values, otherwise multiple imputation approaches can be applied before data analysis. Here we restate the identification assumptions. The independencies in Assumptions 16 should be conditioned additionally on baseline covariates $X$, as listed below.
For notation simplicity, we assume there is no missingness (otherwise we just need to add the missingness indicator in the conditional expectations). In the presence of covariates, one can easily identify and estimate the conditional principal average causal effect (or called the local covariate-specific average treatment effect),
$E\{Y(1)\mid G=cl,X\}$ $=\frac{E\{DSY\mid Z=1,X\}-E\{DSY\mid Z=0,X\}}{E\{DS\mid Z=1,X\}-E\{DS\mid Z=0,X\}},$
$E\{Y(0)\mid G=cl,X\}$ $=\frac{E\{(1-D)SY\mid Z=1,X\}-E\{(1-D)SY\mid Z=0,X\}}{E\{(1-D)S\mid Z=1,X\}-E\{(1-D)S\mid Z=0,X\}}.$
However, when it comes to the population-level average causal effect, the main obstacle for identification is that the proportion of the $cl$ stratum is not identifiable. Indeed, solely with principal ignorability, one can never distinguish $cl$ and $cp$ strata among the treated, nor distinguish $cl$ and $ch$ strata among the controlled using observed data. Therefore, additional assumptions should be made to identify the population-level estimand.
Assumption 7 (Monotonicity on survival)
$S(1)\geq S(0)$ almost surely.
Under monotonicity on survival, the proportions of each principal stratum are all identifiable. Figure 4 illustrates the constitution of observed data in each group. Particularly, the proportion of the $cl$ stratum conditional on covariates is identified as
$P(G=cl\mid X=x)$ $=P(D=0\mid Z=0)E\{S\mid(1-Z)(1-D)\Delta_{S}=1,X=x\}$
$\qquad-P(D=0\mid Z=1)E\{S\mid Z(1-D)\Delta_{S}=1,X=x\}.$
The first term is the conditional proportion of $cl$ and $nl$, and the second term is the conditional proportion of $nl$. Next, the population-level principal average causal effect can be obtained by taking expectation for the conditional principal average causal effect over the covariate distribution on the $cl$ stratum.
Figure 4: Monotonicity on survival. Dark shaded are the observed survivors.
Theorem 4
Under the conditional version of Assumptions 16 and Assumption 7, the principal average causal effect $\tau$ is nonparametrically identifiable,
[…]
where $P(x)$ is the distribution function of $X$ and $\mathcal{X}$ is the support of $X$.
The identification formula motivates a regression estimator for the target estimand. In fact, one can construct weighting estimators using the propensity score and principal score (Feller et al., , 2017; Ding and Lu, , 2017). Then multiply robust estimator can be derived by finding the efficient influence function of the estimand (Wang and Tchetgen Tchetgen, , 2018; Jiang et al., , 2022). When models are correctly specified, the efficient influence function based estimator may enjoy better efficiency than the regression estimator and weighting estimator.
Monotonicity is untestable. Whether monotonicity on survival is reasonable depends on specific scientific questions. There are some variants of the individual-level monotonicity such as stochastic monotonicity (Lee et al., , 2010), which assumes a known relation, also relying untestable prior knowledge, between the potential survivals. To relax the principal ignorability assumption, we can invoke a substitutional (negative control) variable related with potential survivals but not directly related with potential outcomes (Wang et al., , 2017; Deng et al., , 2021). The substitutional variable embeds information of the principal stratum, so the different distribution of mean outcomes can be separated by recovering the information of $G$ using $V$. The data structure would be more complex and may not be easily interpretable for actual scenarios.